OpenStudy (anonymous):

find the value of y that makes the statement true 14-3y=4y

OpenStudy (solomonzelman):

solve for y.

OpenStudy (solomonzelman):

add 3y to both sides,. and divide both sides by 7

OpenStudy (anonymous):

14-3y=4y 14 = 4y + 3y 14 = 7y 14/7 = y y = 14/7